OpenStudy (emilyjones284):

how many grams of sodium hydroxide need to be added to 10 kilograms of water to get a boiling point elevation of the solution of 2 degrees celcius. The Kb for water is 0.512 degrees C/m

OpenStudy (aaronq):

use the formula: \(\Delta T=i*m*K_{\theta}\) where, \(\theta=b\) when finding BP, and \(\theta=f\) when finding freezing point (By that i mean use the appropriate constant.) More on this: http://openstudy.com/study#/updates/521ecb3ae4b0750826e0c362